Sam mixed 9oz of soil containing 12% clay with 5oz of soil with 14% clay. What is the clay content of the mixture?
Answer by macston(5194)

.
T=amount of 12%; F=amount of 14%; x=percent clay (as decimal)
.
0.12(T)+0.14(F) = x(F+T)
(0.12)(9oz)+(0.14)(5oz)=x(14oz)
1.08oz clay + 0.70oz clay = x(14oz)
1.78oz clay=x(14oz)
(1.78 oz/14oz)=x
0.127=x
.
ANSWER: The mixture is 12.7% clay.
.